# KAN-Koopman Based Rapid Detection Of Battery Thermal Anomalies With Diagnostics Guarantees

### v1

@@ Abstract
Early diagnosis of battery thermal anomalies is crucial to ensure safe and reliable battery operation by preventing catastrophic thermal failures. Battery diagnostics primarily rely on battery surface temperature measurements and/or estimation of core temperatures. However, aging-induced changes in the battery model and limited training data remain major challenges for model-based and machine-learning based battery state estimation and diagnostics. To address these issues, we propose a Kolomogorov-Arnold network (KAN) in conjunction with a Koopman-based detection algorithm that leverages the unique advantages of both methods. Firstly, the lightweight KAN provides a model-free estimation of the core temperature to ensure rapid detection of battery thermal anomalies. Secondly, the Koopman operator is learned in real time using the estimated core temperature from KAN and the measured surface temperature of the battery to provide a prediction for diagnostic residual generation. This online learning approach overcomes the challenges of model changes, while the KAN–Koopman integrated structure reduces the dependence on large datasets. Furthermore, we derive analytical conditions that provide diagnostic guarantees on our KAN-Koopman detection scheme. Our simulation results illustrate a significant reduction in detection time with the proposed algorithm compared to the baseline Koopman-only algorithm.
@@ 1 Introduction
Authors proposed a model-free Koopman operator-based detection of internal short circuit- [4] and corrupted charging- [6] induced thermal anomalies in the battery. In addition, the proposed algorithms adopt an online-learning framework that utilizes the limited available data and exhibits an improved adaptability with changes in battery dynamics. However, the methods rely on battery voltage and surface temperature measurements, which can lead to delayed detection of thermal anomalies [7]. Furthermore, while incorporating core temperature information can significantly aid in early detection of thermal anomalies, online core temperature measurement, even with embedded micro-sensors, remains impractical during real-world applications due to its intrusive nature and manufacturing limitations [13]. To address these research gaps, main contributions of this paper are as follows:
@@
The rest of the paper is organized as follows. Section 2 presents a brief overview of KAN network and Koopman operator and Section 3 describes our problem framework. In Section 4, we introduce our proposed KAN-Koopman diagnostic algorithm. The simulation results are presented in Section 5. Finally, Section 6 concludes the paper.
@@ 2.1 Kolomogorov-Arnold network
Let us consider a bounded multi-variate function $f(\boldsymbol{\alpha})$ for $\boldsymbol{\alpha}=(\alpha_{1},\cdots,\alpha_{n})$ that can be expressed with finite additive compositions of $\kappa+1$ times differentiable univariate functions $\varphi_{d,b,a}$ such that $f(\boldsymbol{\alpha})=(\phi_{D}\circ\phi_{D-1}\circ\cdots\circ\phi_{1})\boldsymbol{\alpha}$, where the function matrix $\phi_{d},\,\forall\in\{1,\cdots,D\}$ is defined in (2). Let us also consider a $D$-layer KAN network to approximate this function $f(\boldsymbol{\alpha})$. Then, there exists $\kappa^{th}$ order B-spline functions $\varphi_{d,b,a}^{G}$ learned over a finite grid size $G$ such that the bound on approximation error for the KAN network is:
$\lvert f-(\phi_{D}^{G}\circ\phi_{D-1}^{G}\circ\cdots\circ\phi_{1}^{G})\rvert_{\infty}\leqslant\mathcal{D}G^{-\kappa+1},$ (3)
@@ 4 KAN-Koopman Diagnostic Algorithm
(III) Adaptive residual generation with Koopman model: The offline-trained KAN model fails to account for unforeseen changes in operating conditions. Koopman module, on the other hand, learns the variability in battery dynamics from $T_{2}$ measurements, thereby overcoming the poor adaptability of the KAN network.
Thus, the proposed KAN-Koopman diagnostic algorithm leverages the unique advantages of each method to ensure rapid detection of thermal anomalies with improved adaptability to changes in battery dynamics.
@@ 4.1 KAN-based estimation
$|e_{k}(\textbf{0},0)|_{\infty}\leqslant\mathcal{D}G^{-\kappa+1},\qquad\forall k.$ (14)
@@
Then, using the triangular inequality, (16) becomes
@@
Furthermore, subtracting $\|e_{k}(\textbf{0},0)\|$ on each side of the inequality and choosing $L_{e}:=\max(1,L_{F})$ we prove (15).
@@ 4.3 Diagnostic guarantees
Our previous work in [5] shows that in the presence of anomalies, for a Koopman model that learns from true system measurement, $r_{k}$ will cross the threshold reliably to detect the anomalous input. Hence, we assume that for the hypothetical model described above, there exist a constant $M_{1}\!\in\!\mathbb{R}^{+}$ and practical relaxation constant $\epsilon_{M}\!>\!0$, such that
$\|r_{k}\|\geqslant M_{1}\left[\|\delta_{k}\|+\|\delta_{1_{k}}\|\right]-\epsilon_{M},\quad\forall k.$ (21)
This sensitivity constant $M_{1}\geq 1$ is a design parameter, and a higher value of $M_{1}$ ensures the detection of smaller-valued $\delta_{k},\delta_{1_{k}}$. Furthermore, under no thermal anomalies, the hypothetical residual must remain below the threshold to avoid false detection, i.e. for $\|\delta_{k}\|=\delta_{1_{k}}=0,$ $\|r_{k}\|\!\leqslant\!\epsilon_{M}.$ Next, we will present the two main theorems for our KAN-Koopman framework using these conditions on $r_{k}$.
@@
Consider the proposed KAN-Koopman model (19) learned using the KAN estimate $\mathbb{T}_{1}$ and the hypothetical Koopman model (20) using true measurement $T_{1}$. Then, during nominal battery operation ($\|\delta_{k}\|,\delta_{1_{k}}\!=\!0$), the diagnostic residual $\mathbb{r}_{k}$ is bounded for $\beta_{1},\beta_{2}\in\mathbb{R}^{+}$ and $\varsigma_{k}=\beta_{1}\|C\|+\beta_{2}\|\mathbb{z}_{k}\|$ as:
$\mathbb{r}_{k}\!\leqslant\!r_{k}+\!\left(1+\varsigma_{k}\right)\sqrt{2}\mathcal{D}G^{-\kappa+1},\,\,\,\forall k.$ (22)
@@
Defining $\Delta C:=C$-$\mathbb{C}$, we use the triangular inequality on RHS of (23) and substitute (20) to obtain
@@
Here, $\mathbb{Y}_{b}$ and ${Y}_{b}$ are the delay-embedded measurements for the proposed and hypothetical Koopman models, respectively. Additionally, $\mathbb{z}_{k}$ and $z_{k}$ are the single columns in the Hankel matrices $\mathbb{X}_{b}$ and $X_{b}$, respectively. These matrices $\mathbb{X}_{b}$ and $X_{b}$ contain, respectively, a time-delayed sequence of $\mathbb{y},u$ data and $y,u$ data from previous $k-W_{l}$ observations. Thus, $\Delta z_{k}=\mathbb{z}_{k}-z_{k}$, $\Xi=Y_{b}-\mathbb{Y}_{b}$, and $\zeta=X_{b}-\mathbb{X}_{b}$ contain 0 and $\xi_{j}$ as their entries for $j\in\mathcal{J}\coloneqq\{k,\cdots,k-W_{l}\}$. Let us now define a constant $\beta_{1}>0$ such that $\beta_{1}\|\xi_{k}\|\geqslant W_{l}\max_{j\in\mathcal{J}}(\|\xi_{j}\|)$ and then, we can obtain the bounds as
@@
where $\eta\!=\!X_{b}\zeta^{T}\!+\!\zeta X_{b}^{T}\!-\!\zeta\zeta^{T}$. Now, taking $l_{2}$ norm and using the triangular inequality, we can write $\|\eta\|\!\leqslant\!2\|X_{b}\|\|\zeta\|\!+\!\|\zeta\|^{2}.$ Next, using the boundness of the Hankel matrix $X_{b}$ and $\zeta$ from (27), we can find a constant $\beta_{3}\in\mathbb{R}^{+}$ such that we get $\|\eta\|\leqslant\beta_{3}\|\xi_{k}\|$. Next, multiplying $\left(X_{b}X_{b}^{T}\right)^{-1}X_{b}X_{b}^{T}$ with $(X_{b}X_{b}^{T}-\eta)^{-1}$ from left and merging $X_{b}X_{b}^{T}$, we get
@@
where $\varsigma_{k}=\beta_{1}\|C\|+\beta_{2}\|\mathbb{z}_{k}\|$. Furthermore, in the absence of thermal anomalies, i.e., $\|\delta_{k}\|,\delta_{1_{k}}=0$, (34) becomes
$\mathbb{r}_{k}$ $\leqslant r_{k}+\left(1+\varsigma_{k}\right)\|e_{k}(\textbf{0},0)\|.$ (35)
Additionally, from (14), and using definitions of $l_{2}$ and $l_{\infty}$ norm, we can get $\|e_{k}(\textbf{0},0)\|\leqslant\sqrt{2}|e_{k}(\textbf{0},0)|_{\infty}\leqslant\sqrt{2}\mathcal{D}G^{-\kappa+1}$. Finally, substituting this bound for $\|e_{k}(\textbf{0},0)\|$ in (35), we can obtain (22). This completes the proof. ∎
@@
Theorem 3 (Reliable detection of thermal anomalies).
Consider the proposed KAN-Koopman model (19) learned using the KAN estimate $\mathbb{T}_{1}$ and the hypothetical Koopman model (20) learned using true measurement $T_{1}$. Then, for thermal anomalies $\delta_{k}$ and $\delta_{1_{k}}$ satisfying (37), the bound for residual $\mathbb{r}_{k},\forall k$ can be found for positive constants $M_{1},M_{2}\in\mathbb{R}^{+}$ as:
$\mathbb{r}_{k}$ $\geqslant[M_{1}-M_{2}]\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big]-\epsilon_{M}-\left(1+\varsigma_{k}\right)\!\ \|e_{k}(\textbf{0},0)\|.$ (36)
$\|\delta_{1_{k}}\|+\|\delta_{k}\|\geq\epsilon_{M}+\varsigma_{k}\,\sqrt{2}\mathcal{D}G^{-\kappa+1}.$ (37)
@@
For thermal anomalies $\delta_{k}$ and $\delta_{1_{k}}$ satisfying (37), using the inverse triangular inequality on the RHS of (23), and using the bounds from (27) and (33), (23) becomes
$\mathbb{r}_{k}$ $\geqslant r_{k}-\left(1+\varsigma_{k}\right)\|e_{k}\|.$ (38)
Next, from Lemma 1, we can obtain $\|e_{k}\|\geqslant-\|e_{k}(\textbf{0},0)\|-L_{e}\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big]$. Then, using this bound to substitute $\|e_{k}\|$ and using (21) to substitute $r_{k}$ in (38), we obtain
$\mathbb{r}_{k}$ $\geqslant\left[M_{1}-L_{e}\left(1+\varsigma_{k}\right)\right]\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big]-\epsilon_{M}-\left(1+\varsigma_{k}\right)\|e_{k}(\textbf{0},0)\|.$ (39)
Now, let $M_{2}=L_{e}\left(1+\varsigma_{k}\right)$. Then we can rewrite (39) as (36), which completes the proof. ∎
Theorem 3 implies that if the sensitivity parameter $M_{1}$ is chosen as $M_{1}\geqslant M_{2}$, the diagnostic residual $\mathbb{r}_{k}$ can ensure the reliable detection of thermal anomalies $\delta_{1_{k}}$ and $\delta_{k}$ that satisfy (37), and satisfying (37) indicates thermal anomalies with values greater than the bound of the KAN approximation error $e_{k}(\textbf{0},0)$ and the practical relaxation factor $\epsilon_{M}$. Thus, if $e_{k}(\textbf{0},0)$ is kept small enough with the appropriate network structure (grid $G$ and spline order $\kappa$), the proposed KAN-Koopman algorithm can reliably detect low-magnitude thermal anomalies in battery.

### v2

@@ Abstract
Early diagnosis of battery thermal anomalies is crucial to ensure safe and reliable battery operation by preventing catastrophic thermal failures. Battery diagnostics primarily rely on battery surface temperature measurements and/or estimation of core temperatures. However, aging-induced changes in the battery model and limited training data remain major challenges for model-based and machine-learning based battery state estimation and diagnostics. To address these issues, we propose a Kolomogorov-Arnold network (KAN) in conjunction with a Koopman-based detection algorithm that leverages the unique advantages of both methods. Firstly, the lightweight KAN provides a model-free estimation of the core temperature to ensure rapid detection of battery thermal anomalies. Secondly, the Koopman operator is learned in real time using the estimated core temperature from KAN and the measured surface temperature of the battery to provide the core and surface temperature prediction for diagnostic residual generation. This online learning approach overcomes the challenges of model changes. Furthermore, we derive analytical conditions to obtain diagnostic guarantees on our KAN-Koopman detection scheme. Our simulation results illustrate a significant reduction in detection time with the proposed algorithm compared to the baseline Koopman-only algorithm.
@@ 1 Introduction
Authors proposed a model-free Koopman operator-based detection of internal short circuit- [4] and corrupted charging- [6] induced thermal anomalies in the battery. In addition, the proposed algorithms adopt an online-learning framework that utilizes the limited available data and exhibits an improved adaptability with changes in battery dynamics. However, the methods rely on battery voltage and surface temperature measurements, which can lead to delayed detection of thermal anomalies [7]. Furthermore, while incorporating core temperature information can significantly aid in early detection of thermal anomalies, online core temperature measurement, even with embedded micro-sensors, remains impractical during real-world applications due to its intrusive nature and manufacturing limitations [13, 15]. To address these research gaps, main contributions of this paper are:
@@
The rest of the paper is organized as follows. Section 2 presents a brief overview of the KAN network and the Koopman operator, and Section 3 describes our problem framework. In Section 4, we introduce our proposed KAN-Koopman diagnostic algorithm. The simulation results are presented in Section 5. Finally, Section 6 concludes the paper.
@@ 2.1 Kolomogorov-Arnold network
Let us consider a bounded multi-variate function $f(\boldsymbol{\alpha})$ for $\boldsymbol{\alpha}=(\alpha_{1},\cdots,\alpha_{n})$ that can be expressed with finite additive compositions of $\kappa-1$ times differentiable univariate functions $\varphi_{d,b,a}$ such that $f(\boldsymbol{\alpha})=(\phi_{D}\circ\phi_{D-1}\circ\cdots\circ\phi_{1})\boldsymbol{\alpha}$, where the function matrix $\phi_{d},\,\forall\in\{1,\cdots,D\}$ is defined in (2). Let us also consider a $D$-layer KAN network to approximate this function $f(\boldsymbol{\alpha})$. Then, there exists $\kappa^{th}$ order B-spline functions $\varphi_{d,b,a}^{G}$ learned over a finite grid size $G$ such that the bound on approximation error for the KAN network is:
$\lvert f-(\phi_{D}^{G}\circ\phi_{D-1}^{G}\circ\cdots\circ\phi_{1}^{G})\rvert_{\infty}\leqslant\mathcal{D}G^{-\kappa-1},$ (3)
@@ 4 KAN-Koopman Diagnostic Algorithm
(III) Adaptive residual generation with Koopman model: The offline-trained KAN model fails to account for unforeseen changes in operating conditions. Koopman module, on the other hand, learns the variability in battery dynamics from $T_{2}$ measurements, thereby overcoming the poor adaptability of the KAN network. Thus, the proposed KAN-Koopman diagnostic algorithm leverages the unique advantages of each method to ensure rapid detection of thermal anomalies with improved adaptability to changes in battery dynamics.
@@ 4.1 KAN-based estimation
$|e_{k}(\textbf{0},0)|_{\infty}\leqslant\mathcal{D}G^{-\kappa-1},\qquad\forall k.$ (14)
@@
Then, using the triangle inequality, (16) becomes
@@
Defining $L_{e}:=\max(1,L_{F})$, we obtain
$\|e_{k}\|$ $\leqslant\|e_{k}(\textbf{0},0)\|\,+L_{e}\left(\|\delta_{1_{k}}\|+\,\|\delta_{k}\|\right).$ (19)
subtracting $\|e_{k}(\textbf{0},0)\|$ on each side of the inequality and we prove (15).
@@ 4.3 Diagnostic guarantees
Our previous work in [5] shows that in the presence of anomalies, for a Koopman model that learns from true system measurement, $r_{k}$ will cross the threshold reliably to detect the anomalous input. Hence, we assume that for the hypothetical model described above, there exist a constant $M_{1}\!\in\!\mathbb{R}^{+}$ and a small practical relaxation constant $\epsilon_{M}\!>\!0$, such that
$\|y_{k}-Cz_{k}\|=r_{k}\geqslant M_{1}\left[\|\delta_{k}\|+\|\delta_{1_{k}}\|\right]-\epsilon_{M},\quad\forall k.$ (22)
This sensitivity constant $M_{1}$ is a design parameter, and ensures detection of all attacks that satisfy $\|\delta_{k}\|+\|\delta_{1_{k}}\|\geqslant\frac{\epsilon_{M}}{M_{1}}.$ This bound shows that higher value of sensitivity $M_{1}$ ensures the detection of smaller-valued $\delta_{k},\delta_{1_{k}}$. Furthermore, under no thermal anomalies, the hypothetical residual must remain below the threshold to avoid false detection, i.e.
$r_{k}\!\leqslant\!\epsilon_{M},\qquad\text{when,}\,\,\|\delta_{k}\|,\delta_{1_{k}}=0.$ (23)
Next, we will present the two main theorems for our KAN-Koopman framework using these conditions on $r_{k}$.
@@
Consider the proposed KAN-Koopman model (20) learned using the KAN estimate $\mathbb{T}_{1}$ and the hypothetical Koopman model (21) using true measurement $T_{1}$. Then, during nominal battery operation ($\|\delta_{k}\|,\delta_{1_{k}}\!=\!0$), the diagnostic residual $\mathbb{r}_{k}$ is bounded for $\beta_{1},\beta_{2},\varsigma\in\mathbb{R}^{+}$ as:
$\mathbb{r}_{k}\leqslant\epsilon_{M}+\sqrt{2}\,\varsigma\,\mathcal{D}G^{-\kappa-1},\,\,\,\forall k.$ (24)
@@
Defining $\Delta C:=\mathbb{C}-C$, we use the triangle inequality on RHS of (25) and substitute (21) to obtain
@@
Here, $\mathbb{Y}_{b}$ and ${Y}_{b}$ are the delay-embedded measurements for the proposed and hypothetical Koopman models, respectively. Additionally, $\mathbb{z}_{k}$ and $z_{k}$ are the single columns in the Hankel matrices $\mathbb{X}_{b}$ and $X_{b}$, respectively. These matrices $\mathbb{X}_{b}$ and $X_{b}$ contain, respectively, a time-delayed sequence of $\mathbb{y},u$ data and $y,u$ data from previous $k-W_{l}$ observations. Thus, $\Delta z_{k}=\mathbb{z}_{k}-z_{k}$, $\Xi=\mathbb{Y}_{b}-Y_{b}$, and $\zeta=\mathbb{X}_{b}-X_{b}$ contain 0 and $\xi_{j}$ as their entries for $j\in\mathcal{J}\coloneqq\{k,\cdots,k-W_{l}\}$. Let us now define a constant $\beta_{1}>0$ such that $\beta_{1}\|\xi_{k}\|\geqslant W_{l}\max_{j\in\mathcal{J}}(\|\xi_{j}\|)$ and then, we can obtain the bounds as
@@
where $\eta\!=\!X_{b}\zeta^{T}\!+\!\zeta X_{b}^{T}\!-\!\zeta\zeta^{T}$. Now, taking $l_{2}$ norm and using the triangle inequality, we can write $\|\eta\|\!\leqslant\!2\|X_{b}\|\|\zeta\|\!+\!\|\zeta\|^{2}.$ Next, using the boundness of the Hankel matrix $X_{b}$ and $\zeta$ from (29), we can find a constant $\beta_{3}\in\mathbb{R}^{+}$ such that we get $\|\eta\|\leqslant\beta_{3}\|\xi_{k}\|$. Next, multiplying $\left(X_{b}X_{b}^{T}\right)^{-1}X_{b}X_{b}^{T}$ with $(X_{b}X_{b}^{T}-\eta)^{-1}$ from left and merging $X_{b}X_{b}^{T}$, we get
@@
where $0\leqslant\varsigma_{k}=\beta_{1}\|C\|+\beta_{2}\|\mathbb{z}_{k}\|$. Now, since the Koopman prediction $Cz$ and $\mathbb{C}\mathbb{z}$ remain bounded over the prediction window of length $W_{p}$, we can find a constant $\varsigma\in\mathbb{R}^{+}$ such that $1\leqslant 1+\varsigma_{k}\leqslant\varsigma<\infty,\,\,\forall k$. Hence, in the absence of thermal anomalies, i.e., $\|\delta_{k}\|,\delta_{1_{k}}=0$, (36) becomes
$\mathbb{r}_{k}$ $\leqslant r_{k}+\varsigma\,\|e_{k}(\textbf{0},0)\|.$ (37)
Additionally, from (14), and using definitions of $l_{2}$ and $l_{\infty}$ norm, we can get
$\|e_{k}(\textbf{0},0)\|\leqslant\sqrt{2}|e_{k}(\textbf{0},0)|_{\infty}\leqslant\sqrt{2}\mathcal{D}G^{-\kappa-1}.$ (38)
Finally, substituting this bound for $\|e_{k}(\textbf{0},0)\|$ and $r_{k}$ (23) in (37), we can obtain (24). This completes the proof. ∎
@@
Theorem 3 (KAN-Koopman guarantees for detection).
Consider the proposed KAN-Koopman model (20) learned using the KAN estimate $\mathbb{T}_{1}$ and the hypothetical Koopman model (21) learned using true measurement $T_{1}$. Then, for thermal anomalies $\delta_{k}$ and $\delta_{1_{k}}$ satisfying,
$\|\delta_{1_{k}}\|+\|\delta_{k}\|\geq\epsilon_{M}+\sqrt{2}\,\mathcal{D}G^{-\kappa-1},$ (39)
the residual KAN-Koopman $\mathbb{r}_{k}$ from (20) has the following lower bound for positive constant $M\in\mathbb{R}^{+}$:
$\mathbb{r}_{k}$ $\geqslant M\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big]-\epsilon_{M},\quad\forall k.$ (40)
@@
Using the definitions provided in Theorem 2, let us first rewrite the terms in (25) as
$\mathbb{r}_{k}=\|(y_{k}-Cz_{k})-(C\Delta\mathbb{z}_{k}+\Delta C\mathbb{z}_{k}-\xi_{k})\|.$ (41)
To prove this theorem, we will first show that
$\|C\Delta\mathbb{z}_{k}+\Delta C\mathbb{z}_{k}-\xi_{k}\|\leqslant\|y_{k}-Cz_{k}\|.$ (42)
and then apply the inverse triangle inequality on (41) as
$\mathbb{r}_{k}$ $\geqslant\|y_{k}-Cz_{k}\|-\|C\Delta\mathbb{z}_{k}+\Delta C\mathbb{z}_{k}-\xi_{k}\|.$ (43)
To achieve this goal, we will first use the triangle inequality on the second term to write:
$\|C\Delta\mathbb{z}_{k}+\Delta C\mathbb{z}_{k}-\xi_{k}\|\leqslant\|C\|\,\|\Delta\mathbb{z}_{k}\|+\|\Delta C\|\,\|\mathbb{z}_{k}\|+\|\xi_{k}\|.$ (44)
Next, using the bounds from (29) and (35), the above inequality becomes $\|C\Delta\mathbb{z}_{k}+\Delta C\mathbb{z}_{k}-\xi_{k}\|\leqslant\varsigma\,\|e_{k}\|.$ Subsequently, we first use (19) from Lemma 1 and then (38) from Theorem 2 to this inequality to yield
$\|C\Delta\mathbb{z}_{k}+\Delta C\mathbb{z}_{k}-\xi_{k}\|$ $\leqslant\sqrt{2}\,\varsigma\,\mathcal{D}G^{-\kappa-1}+\varsigma\,L_{e}\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big].$ (45)
Now, for thermal anomalies $\delta_{k}$ and $\delta_{1_{k}}$ satisfying (39),
$\sqrt{2}\,\varsigma\mathcal{D}G^{-\kappa-1}$ $\leqslant\varsigma\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big]-\epsilon_{M}\varsigma,$
$\leqslant\varsigma\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big]-\epsilon_{M},\,\text{(as $\varsigma\geqslant 1$).}$ (46)
Substituting the above inequality in (45) and defining $M_{2}=\varsigma(1+L_{e})$, yields
$\|C\Delta\mathbb{z}_{k}+\Delta C\mathbb{z}_{k}-\xi_{k}\|$ $\leqslant M_{2}\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big]-\epsilon_{M},$ (47)
$\leqslant M_{2}\Big[\|\delta_{1_{k}}\|+\|\delta_{k}\|\Big],$ (48)
as $\epsilon_{M}>0$. Now, we can choose the sensitivity parameter in (22) as $M_{1}>M_{2}$ to claim that $\|y_{k}-Cz_{k}\|=r_{k}\geqslant M_{2}\left[\|\delta_{k}\|+\|\delta_{1_{k}}\|\right]-\epsilon_{M}$. Combining this inequality and (47), we obtain the desired inequality defined in (42). Lastly, using (22) and (48) in the two terms of the inequality (43) respectively, we obtain (40) by defining $M=M_{1}-M_{2}$. This completes the proof. ∎
Theorem 3 implies that if the sensitivity parameter $M_{1}$ is chosen as $M_{1}\geqslant M_{2}$, the diagnostic residual $\mathbb{r}_{k}$ can ensure the reliable detection of thermal anomalies $\delta_{1_{k}}$ and $\delta_{k}$ that satisfy (39), where (39) indicates sensitivity of the method to thermal anomalies with magnitude higher than the bound of the KAN approximation error $e_{k}(\textbf{0},0)$ and the practical relaxation factor $\epsilon_{M}$. Thus, if the bound for $e_{k}(\textbf{0},0)$, i.e., $\sqrt{2}\mathcal{D}G^{-\kappa-1}$ is kept small enough with the appropriate network structure (grid $G$ and spline order $\kappa$), the proposed KAN-Koopman algorithm can reliably detect low-magnitude thermal anomalies in battery.